Circular No. 8798

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2007G, 2007H, 2007I, 2007J, 2007K, 2007L, 2007M Seven apparent supernovae have been discovered on unfiltered CCD

frames: 2007G by C. C. Jin, Y. Cao, F.-Y. Bian, Y.-Q. Lou, X. F. Wang, Y. Wu, and Y. Yang (THCA Supernova Survey/TNT; cf. IAUC 8648); 2007J independently by J. Newton and T. Puckett (cf. IAUC 8794; 0.40-m reflector at Portal, AZ) and by N. Lee and W. Li (LOSS/KAIT; cf. IAUC 8794; source of tabulated data below); and the rest via LOSS (N. Joubert, N. Lee, D. R. Madison, R. Mostardi, and W. Li).

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007G Jan. 9.48 0^{h}08^{m}09.^{s}81 +27^{o}00′11′′.4 17.5 0′′.5 W, 3′′.2 S
2007H Jan. 10.36 8 35 01.98 − 8 20 15.3 17.7 3′′.1 W, 2′′.5 S
2007I Jan. 14.44 11 59 13.15 − 1 36 18.9 18.0 0′′.8 E, 0′′.8 S
2007J Jan. 15.21 2 18 51.70 +33 43 43.3 18.2 5′′.4 E, 15′′.3 N
2007K Jan. 16.40 9 14 26.21 +36 06 52.4 17.2 0′′.3 E, 8′′.1 N
2007L Jan. 18.10 0 43 52.20 + 0 47 35.6 16.8 5′′.1 E, 31′′.3 S
2007M Jan. 14.47 12 04 34.17 +20 11 30.2 19.2 23′′ E, 48′′ S

Additional magnitudes for 2007G in UGC 68: 2006 Oct. 3.55 UT, [18.5 (TNT); Nov. 16.16, [19.0 (KAIT); Dec. 7.15, 17.3 (KAIT); 12.51, 17.5 (TNT); 21.14, 17.3 (KAIT); 2007 Jan. 1.16, 17.4 (KAIT). The tabulated offset for 2007G is with respect to the center of the eastern component in the double-nucleus galaxy UGC 68 (PGC 167). W. Li provides position end figures 09^{s}.78, 11′′.2 for 2007G (and offset 0′′.5 east, 1′′.5 south of the western component of the nucleus of UGC 68). Additional magnitudes for 2007H: 2006 Dec. 19.39, [19.0; 2007 Jan. 14.37, 17.9. Additional magnitudes for 2007I: 2006 Dec. 7.46, [19.2; 2007 Jan. 15.49, 17.6. Additional magnitudes for 2007J in UGC 1778: 2006 Dec. 6.22, [19.5 (KAIT); 18, [19.9 (Puckett); 2007 Jan. 16.14, 17.3 (Puckett); 17.16, 17.3 (Puckett). Additional magnitudes for 2007K in MCG +06-20-50: 2006 Dec. 3.39, [19.5; 2007 Jan. 18.32, 17.1. Additional KAIT magnitudes for 2007L in UGC 466: Jan. 8.11, [18.3; 19.10, 16.8. Additional magnitudes for 2007M in NGC 4076: 2006 May 12.19, [19.5; Dec. 24.51, 18.6; 2007 Jan. 15.51, 19.3. Spectroscopic supernova identifications (and published references): SNe 2007G and 2007L, type-II (CBET 815); 2007H, type-Ia (discovered about a week past maximum; cf. CBET 811); 2007I, type-Ic (discovered ∼ 2 weeks past maximum; cf. CBET 808); 2007J, type-Ia (discovered ≈ 2 months past maximum; cf. CBET 817); 2007K, type-IIn (CBET 813); 2007M, type-Ia (roughly 2 months past maximum on Jan. 22; cf. CBET 816).

2007 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green